Suppose Komla throws a die repeatedly until he gets a six. What is the probability
that he needs to throw more than 10 times to get a six, to 3 decimal places?
1
Expert's answer
2020-06-01T19:32:31-0400

The probability of throwing six in one throw is "\frac{1}{6}", hence the probability of not throwing six in one throw is "1-\frac{1}{6}=\frac{5}{6}."

He needs to throw more than 10 times if and only if he has not thrown six in the first 10 throws.

Since all throws are independent, then the probability of not throwing six in the first 10 throws is "(\frac{5}{6})^{10}\approx 0.162."

Answer: 0.162.